# AL-Bench: A Benchmark for Automatic Logging

### v1

@@ Abstract.
Logging, the practice of inserting log statements into source code, is critical for improving software reliability. Recently, language model-based techniques have been developed to automate log generation based on input code. Although these methods demonstrate promising results in isolated evaluations, their effectiveness diminishes when applied to ad-hoc low-quality data and code similarity-based evaluation methods. We consider a comprehensive evaluation benchmark should include (1) a high-quality, diverse, and large-scale dataset, (2) an assessment of the compilability of the code with inserted log statements, and (3) a runtime log-oriented evaluation method.
To this end, this paper introduces AL-Bench, a comprehensive benchmark designed specifically for automatic logging tools. AL-Bench includes a high-quality, diverse dataset collected from 10 widely recognized projects with varying logging requirements and introduces a novel dynamic evaluation approach. Different from the evaluation in existing logging papers, AL-Bench assesses both the compilability of the code with inserted log statements and the quality of the logs generated by them during runtime, which we believe can better reflect the effectiveness of logging techniques in practice. AL-Bench reveals significant limitations in the state-of-the-art tools The codes with log statements generated by the state-of-the-art tools fail to compile in 20.1%-83.6% cases. In addition, even the best-performing tool did not achieve high similarity between the runtime logs produced by the generated log statements and the ground-truth log statements, demonstrating a 0.213 cosine similarity. The results reveal substantial opportunities to further enhance the development of automatic logging tools.
@@
Numerous logging tools have been developed to assist software developers by automatically suggesting log statements based on provided code snippets (Mastropaolo et al., 2022; Liu et al., 2021; Xu et al., 2024; Xie et al., 2023). As illustrated in Figure 1, automatic logging typically includes three steps: (1) determining the position, (2) deciding the verbosity level, and (3) specifying the message to be recorded. Leveraging the advanced text generation capabilities of large language models (LLMs) (Floridi and Chiriatti, 2020; Raffel et al., 2020), LANCE is the first end-to-end tool to integrate positioning, level selection, and message generation. UniLog (Xu et al., 2024) employed a warm-up and in-context learning (ICL) strategy to enhance performance. FastLog (Xie et al., 2023) improved the generation efficiency while maintaining precision. LEONID (Mastropaolo et al., 2023), based on LANCE, combined with deep learning and information retrieval technologies to enhance performance. SCLogger (Li et al., 2024) adapted static analysis to extend the context for the code snippet. These studies typically evaluate performance using ad-hoc data splitting from the entire dataset, focusing on metrics such as the accuracy of log statement components, including position, level, and message. Machine translation metrics, such as BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004), are also employed to evaluate the quality of generated log messages. Although these evaluations offer valuable insights into logging tool performance, the use of low-quality data and incomplete assessments undermines the reliability of the results.
@@
Second, the current evaluation method does not verify whether the generated log statements are compilable. Generating the compilable log statements is the first requirement when applying automatic logging tools in practice. To relieve developers from the heavy effort required to design and maintain log statements (Chen and Jiang, 2019, 2017), the basic requirement is to ensure that our tool can be seamlessly integrated into the DevOps process without introducing additional errors that require extra debugging effort from developers. Current evaluation methods merely focus on whether each component of log statements (i.e., position, verbosity level, message) matches the ground truth but cannot assess whether the generated log statements might introduce compilation errors such as wrong code format or using undefined variables. Evaluating the compilability of predicted log statements reflects the effectiveness and reliability of logging tools, which are essential for practical use.
Third, the evaluation method cannot evaluate the quality of runtime logs generated by the predicted log statements. Current evaluation methods assess the performance of tools based on the correctness of individual components of log statements. However, current metrics struggle to accurately reflect the quality of runtime logs in real execution environment. Even a slight shift of log statement can lead to the miss of essential runtime details, such as a several-lines shift from the ground truth or a mismatch in verbosity level. For instance, a minor difference in verbosity levels (e.g., debug vs. info) can cause critical information missed due to log level threshold settings in the source code. Therefore, we need to evaluate the quality of log statements in a real execution environment, with the goal of obtaining appropriate logs in specific scenarios, rather than merely focusing on the correctness of individual components or relying on statistical metrics to reflect the performance of logging tools.
@@
    We collected a high-quality, diverse, and large-scale dataset comprising 42,224 instances from 10 popular, high-quality GitHub projects (GitHub, [n. d.]), spanning various domains with differing logging requirements.
@@ 2.1. Log Statement Generation
Logging, the process of generating informative log messages with appropriate verbosity levels at strategically placed locations within code, has long been recognized as a critical challenge in software engineering (He et al., 2021; Chen and Jiang, 2019, 2021; Chen et al., 2021). Over the years, substantial research efforts have aimed to support developers in crafting more effective logging statements, which in turn enhance software maintenance and testing (Jia et al., 2018; Zhao et al., 2017; Zhu et al., 2015). Early studies in this domain often addressed isolated subproblems, typically operating under stringent assumptions that limit the applicability of their findings in real-world scenarios. For example, Li et al. (Li et al., 2021) proposed DeepLV to predict the appropriate logging level by taking surrounding code features into a neural network. Liu et al. (Liu et al., 2022) proposed Tell to further adapted flow graphs to help the suggestions of verbosity levels. Zhu et al. (Zhu et al., 2015) proposed LogAdvisor and Yao et al. (Yao et al., 2019) proposed Log4Perf to assist developers add new log statements in a specific position. Ding et al. proposed LoGenText (Ding et al., 2022) and LoGenText-Plus (Ding et al., 2023) to advise developers what should be logged, and Liu et al. (Liu et al., 2021) proposed tools for deciding which variables should be logged. However, none of them can generate a complete log statement.
@@ 2.2.2. Evaluation Method
First, it cannot assess the compilability of predicted log statements. Compilability is a fundamental requirement for generating appropriate logs. This capability ensures that log statements could be applied into the source code without introducing compilation errors, avoiding wasting developers efforts to debug for the predicted log statements. However, the current evaluation methodology, which focuses solely on comparing the individual components of log statements, cannot identify predicted log statements that use undefined variables or contain incorrect code syntax which could introduce compilation errors. As shown in Figure 4, the predicted log statement is injected into an unreachable position. Applying static evaluation would fail to recognize that this prediction could cause compilation errors. Instead of classifying it as a completely incorrect prediction, the evaluation would count it as a level match instance and assign it a BLEU score of 55.03, contributing to the statistical value. This could introduce significant bias when evaluating the performance, appearing favorable in metrics but performing poorly in actual usage.
@@
Second, it cannot assess the runtime logs generated by predicted log statements. The goal of writing log statements is to obtain the appropriate logs during software execution, so that the quality of log statement is ultimately decided by the quality of runtime logs. Although comparing the correctness of log statement components provides insights into assessing the quality of logs, it can introduce bias when using statistical metrics that focus solely on log statements to evaluate the performance of logging tools. A simple example is a mismatched verbosity level between the predicted log statement and the ground truth. In a case where only the verbosity level is mismatched while the other components are exactly matched, this instance would receive a high score in static evaluation. However, in a real-world scenario, for important error logs, if the predicted level is below the threshold, the logs will not be generated, regardless of how critical the message is. Conversely, for debug logs, if the predicted level exceeds the threshold, these logs could be printed in the production environment, leading to issues such as sensitive information leakage and increased storage costs. These two situations demonstrate that even a shift in verbosity level can turn a log statement into a complete bad case. A more complicated case is presented in Figure 5. Under static evaluation, this prediction with only shift in two lines in position, also could be rewarded highly in statistic metrics. However, it is completely incorrect: placing the log statement at line 2 risks a NullPointerException if the timer object is null when timer.get() is called. This would cause the program to crash during execution, meaning that not only would the logs fail to be generated, but the entire program could terminate unexpectedly. In real-world scenarios, this kind of error is critical and goes unnoticed in static evaluations. While static evaluation rewards this prediction for structural correctness, it fails to account for the actual behavior of the code when executed. Therefore, this example reinforces the need for more comprehensive evaluation methods that go beyond code-level metrics and consider the real-world implications of log statements.
@@
In a conclusion, a log statement’s quality should be determined by its ability to avoid introducing extral errors and generate meaningful, contextually appropriate logs.
@@ 3.1.2. Dataset Construction
In addition to emphasizing data quality, we also addressed the potential risk of data contamination. Since all our data were extracted from public GitHub repositories, which may have been used for training pre-trained models, we implemented precautions to minimize this risk. Specifically, we collected the latest version of each project to reduce the likelihood of it being included in any model’s pre-training data. Furthermore, we wrapped the code snippets in a common class, named A, and standardized the formatting using Google-Java-Format (Google, [n. d.]). This approach altered the format of the code to prevent pre-trained models from recognizing the same information and structure. These strategies have been demonstrated as effective in recent studies (Xu et al., 2024; Tian et al., 2024) Although we adopted effective strategies, data contamination cannot be entirely avoided (Cao et al., 2024). However, our methods minimize this risk and have been proven effective in previous work, providing a solid foundation for analysis and evaluation. In the future, we plan to regularly update our dataset to ensure that the evaluation data remains current. After completing the necessary appeal actions, we finalized our static evaluation dataset.
Table 1. Details of AL-Bench datasets
@@ 3.1.3. Metrics
Metric 1: Level Accuracy (LA): Level Accuracy focuses on log level values, which indicate the importance of the message and are a crucial part of log statements. Some common levels, like ”info,” refer to the normal information of runtime behavior. ”Warning” indicates potential problems that might not immediately cause a disruption but could lead to future issues if not resolved. ”Error” refers to runtime anomalies or issues that need to be addressed. Each level refers to different meanings so that for Level Accuracy we strictly compared the level of predictions and source code and using the exactly matched level number $L_{c}$ divided by the total number of log statements $N_{a}$ to get the value of this metric: $LA=\frac{L_{c}}{N_{a}}$.
@@ 3.2.3. Metrics
Metric 8: False Positive Log Generation Rate (FPLR): This metric measures the proportion of predicted log statements that generate logs during unit test execution when the ground truth log statements would not have produced any logs. It helps assess whether the predicted log statements introduce unnecessary or redundant logs in scenarios where no log should be generated. The number of false positive instances is recorded as $FP$, and the total number of predictions is $P$. he metric is calculated as: $FRLR=\frac{FP}{P}$.
@@ 4. EXPERIMENTS
4.1. Evaluating Automatic Logging Tools
Automatic logging is a hot topic, leading to the development of many tools for determining specific parts of log statements and end-to-end automatic logging tools in recent years. In this paper, we focus on end-to-end logging tools. We reached out to the authors of popular end-to-end automatic logging tools for assistance in rebuilding these tools. However, due to security policies, SCLogger (Li et al., 2024) is still under construction. We ultimately selected four methods for evaluation: LANCE (Mastropaolo et al., 2022), LEONID (Mastropaolo et al., 2023), FastLog (Xie et al., 2023), UniLog (Xu et al., 2024). We detailed the method in the following.
@@ 4.1. Evaluating Automatic Logging Tools
Table 2. Static evaluation on the complete logging task was conducted. The values in the ‘Original’ lines indicate the reported results from the respective methods’ own evaluations. The best performance evaluated by AL-Bench across different metrics is highlighted.
@@ 4.2. RQ1: How well are logging tools at predicting log statements in static evaluation?
After dividing the dataset, we obtained two groups: a longer dataset containing 9,732 instances and a shorter dataset with 32,492 instances. The evaluation results in Table 3 show a clear difference in tool performance based on the length of the instances. Notably, LANCE and LEONID struggled with instances longer than 512 tokens, failing to generate syntax-correct code and, in some cases, producing incomplete code. This explains why their scores for these cases are reported as zero, highlighting the input length limitations of both tools and their inability to handle longer, more complex instances effectively. $UniLog_{cl}$ and FastLog show a considerable drop in PA and DVA when handling longer data, indicating that they struggle to predict log positions and select dynamic variables in more complex instances. $UniLog_{ds}$also shows a dramatic drop in position accuracy, but only a slight decrease in the ability to decide dynamic variables for recording. However, with more sufficient information and a more powerful backbone model, it demonstrates significantly better performance in other metrics. This indicates that with a more powerful backbone, UniLog can better understand the context with longer inputs. Although $UniLog_{d}s$ is equipped with a more powerful backbone, it still struggles to choose the appropriate position for the log statement. As widely demonstrated that the LLMs are not good at counting numbers (Ahn et al., 2024), we might change the output format by directly adding log statement into the source code rather than generate the exactly line number. Furthermore, by simply analyzing the control flow graph of the code, we might be able to exclude positions where logging is not feasible to leave less choices.
@@
Table 4. Compilation Failure Rates across methods
@@
Table 5. Compilation Failure Reasons Analysis

### v2

@@ Abstract.
Logging, the practice of inserting log statements into source code, is critical for improving software reliability. Recently, language model-based techniques have been developed to automate log statement generation based on input code. These tools show promising results in their own evaluation. However, current evaluation practices in log statement generation face significant challenges. The lack of a unified, large-scale dataset forces studies to rely on ad-hoc data, hindering consistency and reproducibility. Additionally, assessments based solely on metrics like code similarity fail to reflect real-world effectiveness. These limitations underscore the need for a comprehensive public benchmark to standardize evaluation.
This paper introduces AL-Bench, a comprehensive benchmark designed specifically for automatic logging tools. AL-Bench includes a high-quality, diverse dataset collected from 10 widely recognized projects with varying logging requirements and introduces a novel dynamic evaluation approach. Different from the existing evaluations that focus only on components of log statements like code similarity, AL-Bench assesses both the compilability of the code with inserted log statements and the effectiveness of the logs generated by them during runtime, which we believe can better reflect the effectiveness of logging techniques in practice. AL-Bench reveals significant limitations in the state-of-the-art tools. The codes with log statements generated by the state-of-the-art tools fail to compile in 20.1%-83.6% cases. In addition, even the best-performing tool only achieves 0.213 cosine similarity between the runtime logs produced by the generated log statements and the ground-truth log statements. The results reveal substantial opportunities to further enhance the development of automatic logging tools.
@@
Numerous logging tools have been developed to assist software developers by automatically suggesting log statements based on provided code snippets (Mastropaolo et al., 2022; Liu et al., 2021; Xu et al., 2024; Xie et al., 2023). As illustrated in Figure 1, automatic logging typically includes three steps: (1) determining the position, (2) deciding the verbosity level, and (3) specifying the message to be recorded. Leveraging the advanced text generation capabilities of large language models (LLMs) (Floridi and Chiriatti, 2020; Raffel et al., 2020), LANCE (Mastropaolo et al., 2022) is the first end-to-end tool to integrate positioning, level selection, and message generation. UniLog (Xu et al., 2024) employed a warm-up and in-context learning (ICL) strategy to enhance performance. FastLog (Xie et al., 2023) improved the generation efficiency while maintaining precision. LEONID (Mastropaolo et al., 2023), based on LANCE, combined with deep learning and information retrieval technologies to enhance performance. SCLogger (Li et al., 2024) adapted static analysis to extend the context for the code snippet. These studies typically evaluate performance using ad-hoc data splitting from the entire dataset, focusing on metrics such as the accuracy of log statement components, including position, level, and message. Machine translation metrics, such as BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004), are also employed to evaluate the quality of generated log messages. Although these evaluations offer valuable insights into logging tool performance, the use of low-quality data and incomplete assessments undermines the reliability of the results.
@@
Second, the current evaluation method does not verify whether the generated log statements are compilable. Generating the compilable log statements is the first requirement when applying automatic logging tools in practice. To relieve developers from the heavy effort required to design and maintain log statements (Chen and Jiang, 2019, 2017), the basic requirement is to ensure that our tool can be seamlessly integrated into the DevOps process without introducing additional errors that require extra debugging effort from developers. Current evaluation methods merely focus on whether each component of log statements (i.e., position, verbosity level, message) matches the ground truth but cannot assess whether the generated log statements might introduce compilation errors such as wrong code format or use undefined variables. Evaluating the compilability of predicted log statements reflects the effectiveness and reliability of logging tools, which are essential for practical use.
Third, the evaluation method cannot evaluate the quality of runtime logs generated by the predicted log statements. Current evaluation methods assess the performance of tools based on the correctness of individual components of log statements. However, current metrics struggle to accurately reflect the quality of runtime logs in the real execution environment. Even a slight shift of log statement can lead to the miss of essential runtime details, such as a several-line shift from the ground truth or a mismatch in verbosity level. For instance, a minor difference in verbosity levels (e.g., debug vs. info) can cause critical information missed due to log level threshold settings in the source code. Therefore, we need to evaluate the quality of log statements in a real execution environment, with the goal of obtaining appropriate logs in specific scenarios, rather than merely focusing on the correctness of individual components or relying on statistical metrics to reflect the performance of logging tools.
@@
    We collected a high-quality, diverse, and large-scale dataset comprising 42,224 instances from 10 popular, high-quality GitHub (GitHub, [n. d.]) projects, spanning various domains with differing logging requirements.
@@ 2.1. Log Statement Generation
Logging, the process of generating informative log messages with appropriate verbosity levels at strategically placed locations within code, has long been recognized as a critical challenge in software engineering (He et al., 2021; Chen and Jiang, 2019, 2021; Chen et al., 2021). Over the years, substantial research efforts have aimed to support developers in crafting more effective logging statements, which in turn enhance software maintenance and testing (Jia et al., 2018; Zhao et al., 2017; Zhu et al., 2015). Early studies in this domain often addressed isolated subproblems, typically operating under stringent assumptions that limit the applicability of their findings in real-world scenarios. For example, Li et al. (Li et al., 2021) proposed DeepLV to predict the appropriate logging level by taking surrounding code features into a neural network. Liu et al. (Liu et al., 2022) proposed Tell to further adapted flow graphs to help the suggestions of verbosity levels. Zhu et al. (Zhu et al., 2015) proposed LogAdvisor and Yao et al. (Yao et al., 2019) proposed Log4Perf to assist developers to add new log statements in a specific position. Ding et al. proposed LoGenText (Ding et al., 2022) and LoGenText-Plus (Ding et al., 2023) to advise developers what should be logged, and Liu et al. (Liu et al., 2021) proposed tools for deciding which variables should be logged. However, none of them can generate a complete log statement.
@@ 2.2.2. Evaluation Method
First, it cannot assess the compilability of predicted log statements. Compilability is a fundamental requirement for generating appropriate logs. This capability ensures that log statements can be applied to the source code without introducing compilation errors, avoiding wasting developers’ efforts to debug for the predicted log statements. However, the current evaluation methodology, which focuses solely on comparing the individual components of log statements, cannot identify predicted log statements that use undefined variables or contain incorrect code syntax which could introduce compilation errors. As shown in Figure 4, the predicted log statement is injected into an unreachable position. Applying static evaluation would fail to recognize that this prediction could cause compilation errors. Instead of classifying it as a completely incorrect prediction, the evaluation would count it as a level match instance and assign it a BLEU score of 55.03, contributing to the statistical value. This could introduce significant bias when evaluating the performance, appearing favorable in metrics but performing poorly in actual usage.
@@
Second, it cannot assess the runtime logs generated by predicted log statements. The goal of writing log statements is to obtain the appropriate logs during software execution so that the quality of log statement is ultimately decided by the quality of runtime logs. Although comparing the correctness of log statement components provides insights into assessing the quality of logs, it can introduce bias when using statistical metrics that focus solely on log statements to evaluate the performance of logging tools. A simple example is a mismatched verbosity level between the predicted log statement and the ground truth. In a case where only the verbosity level is mismatched while the other components are exactly matched, this instance would receive a high score in static evaluation. However, in a real-world scenario, for important error logs, if the predicted level is below the threshold, the logs will not be generated, regardless of how critical the message is. Conversely, for debug logs, if the predicted level exceeds the threshold, these logs could be printed in the production environment, leading to issues such as sensitive information leakage and increased storage costs. These two situations demonstrate that even a shift in verbosity level can turn a log statement into a completely bad case. A more complicated case is presented in Figure 5. Under static evaluation, this prediction with only a shift in two lines in position, also could be rewarded highly in statistic metrics. However, it is completely incorrect: placing the log statement at line 2 risks a NullPointerException if the timer object is null when timer.get() is called. This would cause the program to crash during execution, meaning that not only would the logs fail to be generated, but the entire program could terminate unexpectedly. In real-world scenarios, this kind of error is critical and goes unnoticed in static evaluations. While static evaluation rewards this prediction for structural correctness, it fails to account for the actual behavior of the code when executed. Therefore, this example reinforces the need for more comprehensive evaluation methods that go beyond code-level metrics and consider the real-world implications of log statements.
@@
In conclusion, a log statement’s quality should be determined by its ability to avoid introducing extra errors and generate meaningful, contextually appropriate logs.
@@ 3.1.2. Dataset Construction
In addition to emphasizing data quality, we also addressed the potential risk of data contamination. Since all our data were extracted from public GitHub repositories, which may have been used for training pre-trained models, we implemented precautions to minimize this risk. Specifically, we collected the latest version of each project to reduce the likelihood of it being included in any model’s pre-training data. Furthermore, we wrapped the code snippets in a common class and standardized the formatting using Google-Java-Format (Google, [n. d.]). This approach altered the format of the code to prevent pre-trained models from recognizing the same information and structure. These strategies have been demonstrated as effective in recent studies (Xu et al., 2024; Tian et al., 2024). Although we adopted effective strategies, data contamination cannot be entirely avoided (Cao et al., 2024). However, our methods minimize this risk and have been proven effective in previous work, providing a solid foundation for analysis and evaluation. In the future, we plan to regularly update our dataset to ensure that the evaluation data remains current. After completing the necessary appeal actions, we finalized our static evaluation dataset.
Table 1. Details of AL-Bench dataset.
@@ 3.1.3. Metrics
Metric 1: Level Accuracy (LA): Level Accuracy focuses on log level values, which indicate the importance of the message and are a crucial part of log statements. Some common levels, like ”info,” refer to the normal information of runtime behavior. ”Warning” indicates potential problems that might not immediately cause disruption but could lead to future issues if not resolved. ”Error” refers to runtime anomalies or issues that need to be addressed. Each level refers to different meanings so for Level Accuracy we strictly compared the level of predictions and source code and used the exactly matched level number $L_{c}$ divided by the total number of log statements $N_{a}$ to get the value of this metric: $LA=\frac{L_{c}}{N_{a}}$.
@@ 3.2.3. Metrics
Metric 8: False Positive Log Generation Rate (FPLR): This metric measures the proportion of predicted log statements that generate logs during unit test execution when the ground truth log statements would not have produced any logs. It helps assess whether the predicted log statements introduce unnecessary or redundant logs in scenarios where no log should be generated. The number of false positive instances is recorded as $FP$, and the total number of predictions is $P$. the metric is calculated as: $FRLR=\frac{FP}{P}$.
@@ 4. EXPERIMENTS
4.1. Automatic Logging Tools
Automatic logging is a hot topic, leading to the development of many tools for determining specific parts of log statements and end-to-end automatic logging tools in recent years. In this paper, we focus on end-to-end logging tools. We reached out to the authors of popular end-to-end automatic logging tools for assistance in rebuilding these tools. Because SCLogger (Li et al., 2024) is still under construction, we ultimately selected four methods for evaluation: LANCE (Mastropaolo et al., 2022), LEONID (Mastropaolo et al., 2023), FastLog (Xie et al., 2023), UniLog (Xu et al., 2024). We detailed the method in the following.
@@ 4.1. Automatic Logging Tools
Table 2. Static evaluation on the complete logging task was conducted. The values in the Original lines indicate the reported results from the respective methods’ own evaluations. The best performance evaluated by AL-Bench across different metrics is highlighted.
@@ 4.2. RQ1: How well are logging tools at predicting log statements in static evaluation?
After dividing the dataset, we obtained two groups: a longer dataset containing 9,732 instances and a shorter dataset with 32,492 instances. The evaluation results in Table 3 show a clear difference in tool performance based on the length of the instances. Notably, LANCE and LEONID struggled with instances longer than 512 tokens, failing to generate syntax-correct code and, in some cases, producing incomplete code. This explains why their scores for these cases are reported as zero, highlighting the input length limitations of both tools and their inability to handle longer, more complex instances effectively. $UniLog_{cl}$ and FastLog show a considerable drop in PA and DVA when handling longer data, indicating that they struggle to predict log positions and select dynamic variables in more complex instances. $UniLog_{ds}$also shows a dramatic drop in position accuracy, but only a slight decrease in the ability to decide dynamic variables for recording. However, with more sufficient information and a more powerful backbone model, it demonstrates significantly better performance in other metrics. This indicates that with a more powerful backbone, UniLog can better understand the context with longer inputs. Although $UniLog_{d}s$ is equipped with a more powerful backbone, it still struggles to choose the appropriate position for the log statement. As widely demonstrated that the LLMs are not good at counting numbers (Ahn et al., 2024), we might change the output format by directly adding log statements into the source code rather than generating the exact line number. Furthermore, by simply analyzing the control flow graph of the code, we might be able to exclude positions where logging is not feasible to leave fewer choices.
@@
Table 4. Compilation Failure Rates Across Methods.
@@
Table 5. Compilation Failure Reasons Analysis.